OpenStudy (anonymous):

Find the derivative algebraically using the difference quotient: q(x) = x^-2 at x=2

OpenStudy (xapproachesinfinity):

is that \[\large q(x)=x^{-2}\]

OpenStudy (anonymous):

yes

OpenStudy (xapproachesinfinity):

the diffrence quotient is : \[\large \frac{f(x)-f(2)}{x-2}\]

OpenStudy (xapproachesinfinity):

that's what you need to find out with limit: \[\Large \lim_{x\to2}\frac{f(x)-f(2)}{x-2}=?\]

OpenStudy (anonymous):

I thought the difference quotient is \[\frac{ f(x+h)-f(x) }{ h } \]

OpenStudy (xapproachesinfinity):

well that's the same thing my friend

OpenStudy (anonymous):

\[I got this far but I can't solve the rest: \frac{ -4-h }{ 4h^2+6h+16 }\]

OpenStudy (xapproachesinfinity):

if you want that way it would be \[\large \lim_{h\to0}\frac{f(2+h)-f(2)}{h}\]

OpenStudy (xapproachesinfinity):

what did you use?

OpenStudy (anonymous):

But it is x^-2 not x^2

OpenStudy (xapproachesinfinity):

i'm giving the definition didn't do anything yet my friend

OpenStudy (anonymous):

oh, okay

OpenStudy (xapproachesinfinity):

you are miss using the definition

OpenStudy (xapproachesinfinity):

try to use the one i just wrote

OpenStudy (anonymous):

using what you gave me I got: \[\frac{ \frac{ 1 }{ (2+h)^2 }-\frac{ 1 }{ 4 } }{ h }\]

OpenStudy (xapproachesinfinity):

okay simplify

OpenStudy (anonymous):

\[\frac{ \frac{ -(4+4h+h^2)+4 }{ (4+4h+h^2)4 } }{ h }\]

OpenStudy (xapproachesinfinity):

go further...

OpenStudy (xapproachesinfinity):

carry some cancelations

OpenStudy (anonymous):

that is why I typed that thing earlier I simplified and got to that point and got stuck

OpenStudy (xapproachesinfinity):

well still need to go further... open perenthesis on top and see what can yu spot

OpenStudy (anonymous):

\[Simplified \to \frac{ -4-h }{ 16+6h+4h^2 }\]

OpenStudy (misty1212):

algebra algebra all is algebra...

OpenStudy (xapproachesinfinity):

o don't think so forgot that part and thing about the one you typed just now

OpenStudy (xapproachesinfinity):

-4-4h-h^2+4 ?

OpenStudy (xapproachesinfinity):

that's the first thing to spot

OpenStudy (xapproachesinfinity):

see what you can cancel

OpenStudy (xapproachesinfinity):

@misty1212 pls let him do this :)

OpenStudy (anonymous):

-4h-h^2

OpenStudy (anonymous):

h(4-h)

OpenStudy (xapproachesinfinity):

not quite h(-4-h)

OpenStudy (xapproachesinfinity):

no cancel h top and bottom

OpenStudy (anonymous):

I've gotten to that part

OpenStudy (xapproachesinfinity):

you got incorrect math in there

OpenStudy (xapproachesinfinity):

\[\large \frac{-4-h}{4(2+h)^2}\] oh i see you got to this point

OpenStudy (xapproachesinfinity):

well you need to bring that limit

OpenStudy (xapproachesinfinity):

this is jut difference quotient we need limit to make it derivative

OpenStudy (xapproachesinfinity):

so take the limit: \[\large \lim_{h\to0}\frac{-4-h}{4(2+h)^2}\]

OpenStudy (anonymous):

When I canceled the h from the top and bottom I got: \[\frac{ -4-h }{ (4+4+h^2)4 }\]

OpenStudy (anonymous):

wait h not h^2

OpenStudy (xapproachesinfinity):

no cancelation anymore do the limit i just wrote i said it appears you got the right thing from the beginning

OpenStudy (xapproachesinfinity):

you just had some errors, some number were not correct in denominator

OpenStudy (anonymous):

so I don't have to cancel anymore?

OpenStudy (xapproachesinfinity):

no further cancelation there is nothing to cancel anymore just do the limit

OpenStudy (anonymous):

oh okay, I see now...So I did have it right earlier I just needed to find the limit?

OpenStudy (xapproachesinfinity):

yes sort of, you got the number in the bottom incorrect but other than that it is fine

OpenStudy (anonymous):

so the limit is -1/4

OpenStudy (xapproachesinfinity):

\[\lim_{h\to0}\frac{-4-h}{4(2+h)^2}\] you did this limit?

OpenStudy (anonymous):

yes

OpenStudy (xapproachesinfinity):

how its -1/4?

OpenStudy (xapproachesinfinity):

oh i'm blind hehe you are correct

OpenStudy (xapproachesinfinity):

forgot to the power 2 lol

OpenStudy (anonymous):

\[\frac{ -4-(0) }{ 4(2+0)^2 }\]

OpenStudy (xapproachesinfinity):

yes its correct :) i was not paying attention enough

OpenStudy (anonymous):

sigh...so did I do it correctly before when I wrote \[\frac{ -4-h }{ 16+6h+4h^2 }\]

OpenStudy (xapproachesinfinity):

well the bottom was 4(2+h)^2=16+16h+4h^2

OpenStudy (xapproachesinfinity):

i said you got a math error and that's fine

OpenStudy (anonymous):

oh I see where I got the error

OpenStudy (xapproachesinfinity):

I'm lagging sorry the website is bad again hehe

OpenStudy (anonymous):

I have it written correctly on paper but when I translated it to the computer I wrote 16 instead of 6

OpenStudy (anonymous):

Okay.. thank you

OpenStudy (xapproachesinfinity):

we can check if our result is good: \[\large f'(x)=-2x^{-3}=-2/x^3\] so \[\large f'(2)=-2/8=-1/4\]

OpenStudy (xapproachesinfinity):

you are welcome!